Could I please get assistance for this problem:
In a bread receipe, the ratio of milk to flour is 4/3. If eight cups of milk are used, how many cups of flour are used?

Answer by ichigo449(30)

4/3 = 8/x or x = 8(3/4) = 24/4 = 6. Have a nice day.